Solve the following system of ODE:

x' - y - y' = -e^t
1
Expert's answer
2020-10-12T18:35:40-0400

"x' - y - y' = -e^t\\

\displaystyle\frac{\mathrm{d}x}{\mathrm{d}t} - y - \frac{\mathrm{d}y}{\mathrm{d}t} = -e^t\\

\textsf{If}\hspace{0.1cm} y = x = e^{mt}, \frac{\mathrm{d}y}{\mathrm{d}t} = \frac{\mathrm{d}x}{\mathrm{d}t} = me^{mt}\\

\implies me^{mt} - e^{mt} - me^{mt}= -e^t\\

-e^{mt} = -e^{t} \implies m = 1\\

m = 1\hspace{0.1cm}\textsf{is the only solution}\\\textsf{to the equation.}\\

\therefore y = x = e^{t} \hspace{0.1cm}\textsf{is a solution to the first}\\\textsf{order linear differential equation.}"